Suppose that an industry is dominated by a single producer, the demand for it's product is Q=300-6p. Suppose further that TC=100+5Q+1/480Q×2 is the cost function of the firm.

What is the equilibrium quantity and price in this market given this information


1
Expert's answer
2021-04-14T06:16:21-0400
"TC=100+5Q+\frac {1}{480}Q^2"

"MC=5+\frac{1}{240}Q"

"Q=300-6p"

"p=50-\frac{1}{6}Q"

"5+\frac{1}{240}Q=50-\frac{1}{6}Q"

"Q=263"

"p=6"